# Graph Neural Network based Log Anomaly Detection and Explanation

### v1

@@ 6. Experiments
  3. (3)
    Node Labels vs. Node Attributes: How important is it to use semantic embedding of log event template as node attributes?
  4. (4)
    Robustness analysis: To what extent is Logs2Graphs robust to contamination in training data?
  5. (5)
    Ability to detect structural anomalies: Can Logs2Graphs better capture structural anomalies and identify structurally equivalent normal instances than other contenders?
  6. (6)
    Explainability Analysis: How understandable are the anomaly detection results delivered by Logs2Graphs?
  7. (7)
@@
  8. (8)
    Runtime analysis: What are the runtime for different methods?
6.1. Experiment Setup
6.1.1. Datasets
@@
  • HDFS (Xu et al., 2009) consists of Hadoop Distributed File System logs obtained by running 200 Amazon EC2 nodes. These logs contain block_id, which can be used to group log events into different groups. Moreover, these logs are manually labeled by Hadoop experts.
@@
  • Hadoop (Lin et al., 2016) was collected from a Hadoop cluster consisting of 46 cores over 5 machines. The ContainerID variable is used to divide log messages into different groups.
@@
  • BGL, Spirit and Thunderbird contain system logs collected from the BlueGene/L (BGL) supercomputing system, Spirit supercomputing system, and Thunderbird supercomputing system located at Sandia National Labs, respectively. For those datasets, each log message was manually inspected by engineers and labelled as normal or anomalous. For BGL, we use all log messages, and group log messages based on the Node variable. For Spirit and Thunderbird, we only use the first 1 million and first 5 million log messages for evaluation, respectively. Furthermore, for these two datasets, the User is used as log identifier to group log messages. However, considering that an ordinary user may generate hundreds of thousands of logs, we regard every 100 consecutive logs of each user as a group. If the number of logs is less than 100, we also consider it as a group.
6.1.2. Baselines
@@
To investigate the performance of $Logs2Graphs$, we compare it with the following seven state-of-the-art log anomaly detection methods: Principal Component Analysis (PCA) (Xu et al., 2009), One-Class SVM (OCSVM) (Schölkopf et al., 2001), Isolation Forest (iForest) (Liu et al., 2008), HBOS (Goldstein and Dengel, 2012), DeepLog (Du et al., 2017), LogAnomaly (Meng et al., 2019) and AutoEncoder (Farzad and Gulliver, 2020), and one state-of-the-art graph level anomaly detection method: GLAM (Zhao et al., 2022).
@@
We choose these methods as baselines because they are often regarded to be representatives of traditional machine learning-based (PCA, OCSCM, IForest, HBOS) and deep learning-based approaches (DeepLog, LogAnomaly and AutoEncoder), respectively. All methods are unsupervised or semi-supervised methods that do not require labelled anomalous samples for training the models.
6.1.3. Evaluation Metrics
@@
The Area Under Receiver Operating Characteristics Curve (AUC ROC) and the Area Under the Precision-Recall Curve (AUC PRC) are widely used to quantify the detection accuracy of anomaly detection. Therefore, we employ both to evaluate and compare the different log anomaly detection methods. AUC PRC is also known as Average Precision (AP). For both AUC ROC and AUC PRC, values closer to 1 indicate better performance.
6.2. Model Implementation and Configuration
@@
Traditional machine learning based approaches—such as PCA, OCSVM, iForest, and HBOS—usually first transform logs into log event count vectors, and then apply traditional anomaly detection techniques to identify anomalies. For these methods, we utilise their open-source implementations provided in PyOD (Zhao et al., 2019). Meanwhile, for deep learning methods DeepLog, LogAnomaly, and AutoEncoder, we use their open-source implementations in Deep-Loglizer (Chen et al., 2021). For these competing methods, we use their default hyperparameter values.
@@
For all deep learning based methods, the experimental design adopted in this study follows a train/validation/test strategy with a distribution of $70\%:5\%:25\%$ for normal instances. Specifically, the model was trained using $70\%$ of normal instances, while $5\%$ of normal instances and an equal number of abnormal instances were employed for validation (i.e., hyperparameter tuning). The remaining $25\%$ of normal instances and the remaining abnormal instances were used for testing. Specifically, Table 2 summarises the hyperparameters involved in OCDiGCN as well as their recommended values.
@@
We implemented and ran all algorithms in Python 3.8 (using PyTorch (Paszke et al., 2019) and PyTorch Geometric (Fey and Lenssen, 2019) libraries when applicable), on a computer with Apple M1 chip 8-core CPU and 16GB unified memory. For reproducibility, all code and datasets will be released on GitHub.
6.3. Comparison to the state of the art
@@
We first compare $Logs2Graphs$ to the state of the art. The results are shown in Table 3, based on which we make the following main observations:
@@
  • In terms of AUC ROC, Logs2Graphs achieves the best performance against its competitors on four out of five datasets. Particularly, Logs2Graphs outperforms the closet competitor on BGL with 9.6% and delivers remarkable results (i.e., an AUC ROC larger than 0.99) on Spirit and Thunderbird. Similar observations can be made for Average Precision.
@@
  • The performance of (not-graph-based) deep learning methods is often inferior to that of Log2Graphs on the more complex datasets, i.e., Hadoop, BGL, Spirit, and Thunderbird, which all contain hundreds or even thousands of log templates. This suggests that LSTM-based models may not be well suited for logs with a large number of log templates. One possible reason is that the test dataset contains many unprecedented log templates, namely log templates that are not present in the training dataset.
@@
Table 3. Anomaly detection accuracy on five benchmark datasets for Logs2Graphs and its eight competitors. AP and RC denote Average Precision and AUC ROC, respectively. HDFS, BGL, and Thunderbird have been downsampled to 10,000 graphs each while maintaining the original anomaly rates. For each method on each dataset, to mitigate potential biases arising from randomness, we conducted ten experimental runs with varying random seeds and report the average values along with standard deviations of AP and RC. Moreover, we highlight the best results with bold and the runner-up with underline.
HDFS Hadoop BGL Spirit Thunderbird
Method
AP
RC
AP
RC
AP
RC
AP
RC
AP
RC
PCA
0.91$\pm$0.03
1.0$\pm$0.00
0.84$\pm$0.00
0.52$\pm$0.00
0.73$\pm$0.01
0.82$\pm$0.00
0.31$\pm$0.00
0.19$\pm$0.00
0.11$\pm$0.00
0.34$\pm$0.01
OCSVM
0.18$\pm$0.01
0.88$\pm$0.01
0.83$\pm$0.00
0.45$\pm$0.00
0.47$\pm$0.00
0.47$\pm$0.01
0.34$\pm$0.00
0.29$\pm$0.00
0.12$\pm$0.00
0.45$\pm$0.01
IForest
0.73$\pm$0.04
0.97$\pm$0.01
0.85$\pm$0.01
0.55$\pm$0.01
0.79$\pm$0.01
0.83$\pm$0.01
0.32$\pm$0.03
0.23$\pm$0.02
0.11$\pm$0.01
0.24$\pm$0.10
HBOS
0.74$\pm$0.04
0.99$\pm$0.00
0.84$\pm$0.00
0.50$\pm$0.00
0.84$\pm$0.02
0.87$\pm$0.03
0.35$\pm$0.00
0.22$\pm$0.00
0.15$\pm$0.01
0.29$\pm$0.05
DeepLog
0.92$\pm$0.07
0.97$\pm$0.04
0.96$\pm$0.00
0.47$\pm$0.00
0.89$\pm$0.00
0.72$\pm$0.00
0.99$\pm$0.00
0.97$\pm$0.00
0.91$\pm$0.01
0.96$\pm$0.00
LogAnomaly
0.89$\pm$0.09
0.95$\pm$0.05
0.96$\pm$0.00
0.47$\pm$0.00
0.89$\pm$0.00
0.72$\pm$0.00
0.99$\pm$0.00
0.97$\pm$0.00
0.90$\pm$0.01
0.96$\pm$0.00
AutoEncoder
0.71$\pm$0.03
0.84$\pm$0.01
0.96$\pm$0.00
0.52$\pm$0.00
0.91$\pm$0.01
0.79$\pm$0.02
0.96$\pm$0.00
0.92$\pm$0.01
0.44$\pm$0.02
0.46$\pm$0.05
GLAM
0.78$\pm$0.08
0.89$\pm$0.04
0.95$\pm$0.00
0.61$\pm$0.00
0.94$\pm$0.02
0.90$\pm$0.03
0.93$\pm$0.00
0.91$\pm$0.00
0.75$\pm$0.02
0.85$\pm$0.01
Logs2Graphs
0.87$\pm$0.04
0.91$\pm$0.02
0.95$\pm$0.00
0.59$\pm$0.00
0.96$\pm$0.01
0.93$\pm$0.01
1.0$\pm$0.00
1.0$\pm$0.00
0.99$\pm$0.00
1.0$\pm$0.00
6.4. Directed vs. undirected graphs
To investigate the practical added value of using directed log graphs as opposed to undirected log graphs, we convert the logs to attributed, undirected, and edge-weighted graphs, and apply GLAM (Zhao et al., 2022), a graph-level anomaly detection method for undirected graphs. We use the same graph construction method as for Logs2Graphs, except that we use undirected edges. Similar to our method, GLAM also couples the graph representation learning and anomaly detection objectives by optimising a single SVDD objective. The key difference with OCDiGCN is that GLAM leverages GIN (Xu et al., 2018), which can only tackle undirected graphs, while OCDiGCN utilises DiGCN (Tong et al., 2020) that is especially designed for directed graphs.
@@
6.5. Node Labels vs. Node Attributes
To investigate the importance of using semantic embedding of log event template as node attributes, we replace the node semantic attributes with one-hot-encoding of node labels (i.e., using an integer to represent a log event template). The performance comparisons in terms of ROC AUC for Logs2Graphs are depicted in Figure 3, which shows that using semantic embedding is always superior to using node labels. Particularly, it can lead to a substantial performance improvement on Hadoop, Spirit and HDFS datasets. The PR AUC results show a similar behaviour and thus are omitted.
Figure 3. The comparative performance analysis of Logs2Graphs, measured by ROC AUC, demonstrating the distinction between utilizing node semantic attributes and node labels.
6.6. Robustness to Contamination
To investigate the robustness of Logs2Graphs when the training dataset is contaminated, we report its performance in terms of ROC AUC under a wide range of contamination levels. Figure 4 shows that the performance of Logs2Graphs decreases with an increase of contamination in the training data. The PR AUC results show a similar behaviour and thus are omitted. Hence, it is important to ensure that the training data contains only normal graphs (or with a very low proportion of anomalies).
Figure 4. ROC results of Logs2Graphs w.r.t. a wide range of contamination levels. Results are averaged over 10 runs. Particularly, HDFS contains only 3% anomalies and thus results at 5% and 10% are not available.
6.7. Ability to Detect Structural Anomalies and Recognise Unseen Normal Instances
To showcase the effectiveness of different neural networks in detecting structural anomalies, we synthetically generate normal and anomalous directed graphs as shown in Figure 5. As Deeplog, LogAnomaly and AutoEncoder require log sequences as inputs, we convert directed graphs into sequences by sequentially presenting the endpoints pair of each edge. Moreover, for GLAM we convert directed graphs into undirected graphs by turning each directed edge into undirected edge.
Figure 5. Synthetic generation of normal (10000) and structurally anomalous (200 each) graphs.
@@ 6.7. Ability to Detect Structural Anomalies and Recognise Unseen Normal Instances
Specifically, the results in Table 4 indicate that Logs2Graphs, Deelog and LogAnomaly can effectively detect structural anomalies while AutoEncoder and GLAM fail in some cases. However, log sequences based methods, namely Deelog, LogAnomaly and AutoEncoder, can lead to high positive rates due to their inability of recognising unseen but structurally equivalent normal instances.
Table 4. ROC AUC Results (higher is better) of detecting structural anomalies and False Positive Rate (lower is better) of recognising unseen normal instances. S1: Reverse Edge Direction; S2: Change Edge Endpoint; S3: Delete Edge; S4: Add Edge; N1: Unseen normal instances.
Case
Deeplog
LogAnomaly
AutoEncoder
GLAM
Ours
S1 (RUC)
1.0
1.0
0.0
0.0
1.0
S2 (RUC)
1.0
1.0
0.50
1.0
1.0
S3 (RUC)
1.0
1.0
1.0
1.0
1.0
S4 (RUC)
1.0
1.0
1.0
1.0
1.0
N1 (FPR)
100%
100%
100%
0%
0%
6.8. Anomaly Explanation
Particularly, Figure 6 provides an example of log anomaly explanation with the HDFS dataset. For each detected anomalous log graph (namely a group of logs), we first quantify the importance of nodes according to the description in Section 5.3. Next, we visualise the anomalous graph by assigning darker shade of red to more important nodes. In this example, the node “WriteBlock(WithException)” contributes the most to the anomaly score of an anomalous log group and thus is highlighted in red.
Figure 6. An example of anomaly explanation with HDFS dataset (the log event templates are simplified for better visualisation).
6.9. Sensitivity Analysis
Figure 7. The effects of the number of layers (top row), the embedding dimensions (middle row) and the proximity parameter (bottom row) on AP (left column) and AUC ROC (right column).

### v3

@@ 6 Experiments
  3. 3.
    Node Labels vs. Node Attributes: How important is it to use semantic embeddings of log events as node attributes?
  4. 4.
    Robustness analysis: To what extent is Logs2Graphs robust to contamination of the training data?
  5. 5.
    Ability to detect structural anomalies: Can Logs2Graphs better capture structural anomalies and identify structurally equivalent normal instances than its contenders?
  6. 6.
    Explainability Analysis: How understandable are the anomaly explanations given by Logs2Graphs?
  7. 7.
@@
  8. 8.
    Runtime analysis: What are the runtimes for the different methods?
6.1 Experiment Setup
6.1.1 Datasets
@@
    HDFS Xu et al. (2009) consists of Hadoop Distributed File System logs obtained by running 200 Amazon EC2 nodes. These logs contain block_id, which can be used to group log events into different groups. Moreover, these logs are manually labeled by Hadoop experts.
  • 2.
@@
    Hadoop Lin et al. (2016) was collected from a Hadoop cluster consisting of 46 cores over 5 machines. The ContainerID variable is used to divide log messages into different groups.
  • 3.
@@
    BGL, Spirit, and Thunderbird contain system logs collected from the BlueGene/L (BGL), Spirit, and Thunderbird supercomputing systems located at Sandia National Labs, respectively. For those datasets, each log message was manually inspected by engineers and labelled as normal or anomalous. For BGL, we use all log messages, and group log messages based on the Node variable. For Spirit and Thunderbird, we only use the first 1 million and first 5 million log messages for evaluation, respectively. Furthermore, for these two datasets, the User is used as log identifier to group log messages. However, considering that an ordinary user may generate hundreds of thousands of logs, we regard every 100 consecutive logs of each user as a group. If the number of logs is less than 100, we also consider it as a group.
6.1.2 Baselines
@@
To investigate the performance of $Logs2Graphs$, we compare it with the following seven log anomaly detection methods: Principal Component Analysis (PCA) Xu et al. (2009), One-Class SVM (OCSVM) Schölkopf et al. (2001), Isolation Forest (iForest) Liu et al. (2008), HBOS Goldstein and Dengel (2012), DeepLog Du et al. (2017), LogAnomaly Meng et al. (2019), and AutoEncoder Farzad and Gulliver (2020), and one state-of-the-art graph level anomaly detection method: GLAM Zhao et al. (2022).
@@
We choose these methods as baselines because they are often regarded to be representatives of traditional machine learning-based (PCA, OCSVM, IForest, HBOS) and deep learning-based approaches (DeepLog, LogAnomaly and AutoEncoder), respectively. All methods are unsupervised or semi-supervised methods that do not require labelled anomalous samples for training the models.
6.1.3 Evaluation Metrics
@@
The Area Under Receiver Operating Characteristics Curve (ROC AUC) and the Area Under the Precision-Recall Curve (PRC AUC) are widely used to quantify the detection accuracy of anomaly detection Aggarwal and Aggarwal (2017). This is mainly because they can provide a single value that summarizes the overall performance of the anomaly detection model across various thresholds. In contrast, other metrics such as Precision, Recall and F1-score depend on choosing a threshold to determine whether an instance is anomalous or normal. Consequently, different thresholds can result in different values. Therefore, we employ ROC AUC and PRC AUC to evaluate and compare the different log anomaly detection methods. PRC AUC is also known as Average Precision (AP). For both ROC AUC and PRC AUC, values closer to 1 indicate better performance.
6.2 Model Implementation and Configuration
@@
Traditional machine learning based approaches—such as PCA, OCSVM, iForest, and HBOS—usually first transform logs into log event count vectors, and then apply traditional anomaly detection techniques to identify anomalies. For these methods, we utilise their open-source implementations provided in PyOD Zhao et al. (2019). Meanwhile, for deep learning methods DeepLog, LogAnomaly, and AutoEncoder, we use their open-source implementations in Deep-Loglizer Chen et al. (2021). For these methods, we use their default hyperparameter values.
@@
For all deep learning based methods, the experimental design adopted in this study follows a train/validation/test strategy with a distribution of $70\%:5\%:25\%$ for normal instances. Specifically, the model was trained using $70\%$ of normal instances, while $5\%$ of normal instances and an equal number of abnormal instances were employed for validation (i.e., hyperparameter tuning). The remaining $25\%$ of normal instances and the remaining abnormal instances were used for testing. Table 2 summarises the hyperparameters involved in OCDiGCN as well as their recommended values.
@@
We implemented and ran all algorithms in Python 3.8 (using PyTorch Paszke et al. (2019) and PyTorch Geometric Fey and Lenssen (2019) libraries when applicable), on a workstation equipped with an Intel i7-11700KF CPU and Nvidia RTX3070 GPU. For reproducibility, all code and datasets are released on GitHub. [^2] https://github.com/ZhongLIFR/Logs2Graph
Table 3: Anomaly detection accuracy on five benchmark datasets for Logs2Graphs and its eight competitors. AP and RC denote Average Precision and ROC AUC, respectively. HDFS, BGL, and Thunderbird have been downsampled to 10,000 graphs each while maintaining the original anomaly rates. For each method on each dataset, to mitigate potential biases arising from randomness, we conducted ten experimental runs with varying random seeds and report the average values along with standard deviations of AP and RC. Moreover, we highlight the best results with bold and the runner-up with underline.
HDFS Hadoop BGL Spirit Thunderbird Method AP RC AP RC AP RC AP RC AP RC PCA 0.91$\pm$0.03 1.0$\pm$0.00 0.84$\pm$0.00 0.52$\pm$0.00 0.73$\pm$0.01 0.82$\pm$0.00 0.31$\pm$0.00 0.19$\pm$0.00 0.11$\pm$0.00 0.34$\pm$0.01 OCSVM 0.18$\pm$0.01 0.88$\pm$0.01 0.83$\pm$0.00 0.45$\pm$0.00 0.47$\pm$0.00 0.47$\pm$0.01 0.34$\pm$0.00 0.29$\pm$0.00 0.12$\pm$0.00 0.45$\pm$0.01 IForest 0.73$\pm$0.04 0.97$\pm$0.01 0.85$\pm$0.01 0.55$\pm$0.01 0.79$\pm$0.01 0.83$\pm$0.01 0.32$\pm$0.03 0.23$\pm$0.02 0.11$\pm$0.01 0.24$\pm$0.10 HBOS 0.74$\pm$0.04 0.99$\pm$0.00 0.84$\pm$0.00 0.50$\pm$0.00 0.84$\pm$0.02 0.87$\pm$0.03 0.35$\pm$0.00 0.22$\pm$0.00 0.15$\pm$0.01 0.29$\pm$0.05 DeepLog 0.92$\pm$0.07 0.97$\pm$0.04 0.96$\pm$0.00 0.47$\pm$0.00 0.89$\pm$0.00 0.72$\pm$0.00 0.99$\pm$0.00 0.97$\pm$0.00 0.91$\pm$0.01 0.96$\pm$0.00 LogAnomaly 0.89$\pm$0.09 0.95$\pm$0.05 0.96$\pm$0.00 0.47$\pm$0.00 0.89$\pm$0.00 0.72$\pm$0.00 0.99$\pm$0.00 0.97$\pm$0.00 0.90$\pm$0.01 0.96$\pm$0.00 AutoEncoder 0.71$\pm$0.03 0.84$\pm$0.01 0.96$\pm$0.00 0.52$\pm$0.00 0.91$\pm$0.01 0.79$\pm$0.02 0.96$\pm$0.00 0.92$\pm$0.01 0.44$\pm$0.02 0.46$\pm$0.05 GLAM 0.78$\pm$0.08 0.89$\pm$0.04 0.95$\pm$0.00 0.61$\pm$0.00 0.94$\pm$0.02 0.90$\pm$0.03 0.93$\pm$0.00 0.91$\pm$0.00 0.75$\pm$0.02 0.85$\pm$0.01 Logs2Graphs 0.87$\pm$0.04 0.91$\pm$0.02 0.95$\pm$0.00 0.59$\pm$0.00 0.96$\pm$0.01 0.93$\pm$0.01 1.0$\pm$0.00 1.0$\pm$0.00 0.99$\pm$0.00 1.0$\pm$0.00
6.3 Comparison to the state of the art
@@
We first compare $Logs2Graphs$ to the state of the art. We have the following main observations according to the results in Table 3:
  • 1.
@@
    In terms of ROC AUC, Logs2Graphs achieves the best performance against its competitors on three out of five datasets. Particularly, Logs2Graphs outperforms the closet competitor on BGL with 9.6% and delivers remarkable results (i.e., an ROC AUC larger than 0.99) on Spirit and Thunderbird. Similar observations can be made for Average Precision.
  • 2.
@@
  • 3.
    The performance of (not-graph-based) deep learning methods is often inferior to that of Log2Graphs on the more complex datasets, i.e., BGL, Spirit, and Thunderbird, which all contain hundreds or even thousands of log templates. This suggests that LSTM-based models may not be well suited for logs with a large number of log templates. One possible reason is that the test dataset contains many unprecedented log templates, namely log templates that are not present in the training dataset.
  • 4.
@@
6.4 Directed vs. undirected graphs
To investigate the practical added value of using directed log graphs as opposed to undirected log graphs, we convert the logs to attributed, undirected, and edge-weighted graphs, and apply GLAM Zhao et al. (2022), a graph-level anomaly detection method for undirected graphs. We use the same graph construction method as for Logs2Graphs, except that we use undirected edges. Similar to our method, GLAM also couples the graph representation learning and anomaly detection objectives by optimising a single SVDD objective. The key difference with OCDiGCN is that GLAM leverages GIN Xu et al. (2018), which can only tackle undirected graphs, while OCDiGCN utilises DiGCN Tong et al. (2020) that is especially designed for directed graphs.
@@
6.5 Node Labels vs. Node Attributes
To investigate the importance of using semantic embeddings of log events as node attributes, we replace the node semantic attributes with one-hot-encoding of node labels (i.e., using an integer to represent a log event). The performance comparisons in terms of ROC AUC for Logs2Graphs are depicted in Figure 3, which shows that using semantic embeddings is always superior to using node labels. Particularly, it can lead to a substantial performance improvement on the Hadoop, Spirit and HDFS datasets. The PRC AUC results show a similar behaviour and thus are omitted.
Figure 3: The comparative performance analysis of Logs2Graphs, measured by ROC AUC, demonstrating the distinction between utilizing node semantic attributes and node labels.
6.6 Robustness to Contamination
To investigate the robustness of Logs2Graphs when the training dataset is contaminated (namely integrating anomalous graphs in training data), we report its performance in terms of ROC AUC under a wide range of contamination levels. Figure 4 shows that the performance of Logs2Graphs decreases with an increase of contamination in the training data. The PRC AUC results show a similar behaviour and thus are omitted. Hence, it is important to ensure that the training data contains only normal graphs (or with a very low proportion of anomalies).
Figure 4: ROC AUC results of Logs2Graphs w.r.t. a wide range of contamination levels. Results are averaged over 10 runs. Particularly, HDFS contains only 3% anomalies and thus results at 5% and 10% are not available.
6.7 Ability to Detect Structural Anomalies and Recognise Unseen Normal Instances
To showcase the effectiveness of different neural networks in detecting structural anomalies, we synthetically generate normal and anomalous directed graphs as shown in Figure 5. As Deeplog, LogAnomaly and AutoEncoder require log sequences as inputs, we convert directed graphs into sequences by sequentially presenting the endpoints pair of each edge. Moreover, for GLAM we convert directed graphs into undirected graphs by turning each directed edge into an undirected edge.
Figure 5: Synthetic generation of normal (10000) and structurally anomalous (200 each) graphs.
@@ 6.7 Ability to Detect Structural Anomalies and Recognise Unseen Normal Instances
The results in Table 4 indicate that Logs2Graphs, Deeplog and LogAnomaly can effectively detect structural anomalies, while AutoEncoder and GLAM fail in some cases. However, log sequences based methods, namely Deeplog, LogAnomaly and AutoEncoder, can lead to high false positive rates due to their inability of recognising unseen but structurally equivalent normal instances.
Table 4: ROC AUC results (higher is better) of detecting structural anomalies and False Positive Rate (lower is better) of recognising unseen normal instances. S1: Reverse Edge Direction; S2: Change Edge Endpoint; S3: Delete Edge; S4: Add Edge; N1: Unseen normal instances.
Case Deeplog LogAnomaly AutoEncoder GLAM Ours
S1 (ROC) 1.0 1.0 0.0 0.0 1.0
S2 (ROC) 1.0 1.0 0.50 1.0 1.0
S3 (ROC) 1.0 1.0 1.0 1.0 1.0
S4 (ROC) 1.0 1.0 1.0 1.0 1.0
N1 (FPR) 100% 100% 100% 0% 0%
6.8 Anomaly Explanation
Figure 6 provides an example of log anomaly explanation with the HDFS dataset. For each detected anomalous log graph (namely a group of logs), we first quantify the importance of nodes according to the description in Section 5.3. Next, we visualise the anomalous graph by assigning darker shade of red to more important nodes. In this example, the node “WriteBlock(WithException)” contributes the most to the anomaly score of an anomalous log group and thus is highlighted in red.
Figure 6: Example of anomaly explanation with HDFS (the log event templates are simplified for better visualisation).
6.9 Sensitivity Analysis
Figure 7: The effects of the number of layers (top row), the embedding dimensions (middle row) and the proximity parameter (bottom row) on AP (left column) and ROC AUC (right column).